Multiply.
(2r-3s)(4r squared +6rs + 9s squared)
Please help!

Answer by apachemexican1(64)

(2r-3s)(4r^2+6rs+9s^2)
use foil to solve this
first multiply
2r*4r^2=8r^5
2r*6rs=12r^2s
2r*9s^2=18rs^2
now you have half the solution
second multiply
-3s*4r^2=-12r^2s
-3s*6rs=-18rs^2
-3s*9s^2=27s^3
now you half the other half of the solution
combine like terms
8r^5+12r^2s-12r^2s+18rs^2-18rs^2+27s^3=
8r^5+27s^3
and that is the solution
good luck!